OpenStudy (anonymous):

Simplify: x squared minus 10x + 9 over x squared + 2x minus 3

OpenStudy (anonymous):

\[x ^{2}-10x+9 \over x ^{2}+2x-3\]

OpenStudy (anonymous):

x^2 - 10x + 9 = (x-1) (x-9) x^2 + 2x - 3 = (x - 1) (x + 3)

OpenStudy (anonymous):

\[\frac{(x-1)(x-9)}{(x-1)(x+3)}\]

jhonyy9 (jhonyy9):

(x-9)(x-1) ---------- (x-1)(x+3)

OpenStudy (anonymous):

so cancels x - 1

OpenStudy (anonymous):

\[\frac{(x-9)}{(x+3)}\]

OpenStudy (anonymous):

Thanks :)